OpenStudy (anonymous):

simplify (x^-3y^5z^0/x^2y^8)^2

OpenStudy (anonymous):

\[\frac{ x^-3y^5z^0 }{ x^2y^8 }^2\]

OpenStudy (anonymous):

()^2

OpenStudy (anonymous):

\[(x ^{-3}y ^{5}z ^{0}/x ^{2}y ^{8})^{2}\] simplify it by dividing term by term like x^-3/x^2 =x^-3-2=? y^5/y^3=y^5-3=? also what is z^0=?

OpenStudy (anonymous):

x^6

OpenStudy (anonymous):

y^-15

OpenStudy (anonymous):

\[\frac{ x ^{a} }{ x ^{b} }=x ^{a-b}\]

OpenStudy (anonymous):

\[\left(\begin{matrix}x^-3y^5z^0 \\ x^2y^8\end{matrix}\right)^2\]

OpenStudy (anonymous):

thats the problem

OpenStudy (anonymous):

i'm new at this

OpenStudy (anonymous):

yeah later on you squared them after simplifying the terms

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i'm so lost with these

OpenStudy (anonymous):

but first simlipfying the division of powers are adding and subtracting on not multiplying .. :D

OpenStudy (anonymous):

so x^1y^-3z^0

OpenStudy (anonymous):

\[\frac{ z }{ x^2y^6 }\]

OpenStudy (anonymous):

\[\frac{ y ^{-c} }{ y ^{b} }=y ^{-c-b}\]

OpenStudy (anonymous):

look i dont understand

OpenStudy (anonymous):

\[\frac{ v ^{-c} }{ y ^{b} }=y ^{-c}*y ^{-b}=y ^{-c-b}\]

OpenStudy (anonymous):

i get the x^-1*2 y^-3*2

OpenStudy (anonymous):

for example only \[\frac{ y ^{-5} }{ y ^{7} }=y ^{-5}*y ^{-7}=y ^{-5-7}=y ^{-12}\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

when you move the denominator to the numerator the sign of its power is change

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\left(\begin{matrix}x^0 \\ x^1y^3\end{matrix}\right)^2\]

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

\[\left(\begin{matrix}z \\ x^1y^3\end{matrix}\right)^2\]

OpenStudy (anonymous):

that top is z

OpenStudy (anonymous):

hm the idea is to move all the denominator upward to the numerator first then do add or subtract

OpenStudy (anonymous):

so x^-3-2y^5-8z^0

OpenStudy (anonymous):

yes correct and z^0=1 right?

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

so x^-1y-3z^1

OpenStudy (anonymous):

any number or letter raised to the power of zero is =1

OpenStudy (anonymous):

oh right i have been working on homework for the last 5 hours my mind is fried =)

OpenStudy (anonymous):

hmm what is -2-3=? 5-8=?

OpenStudy (anonymous):

your 5-8= -3 is correct :D

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yesss :D

OpenStudy (anonymous):

yess =)

OpenStudy (anonymous):

so whats the answer now?

OpenStudy (anonymous):

so x^-5y^-3x^1

OpenStudy (anonymous):

than that times 2

OpenStudy (anonymous):

so x^-5*2=-10 y^-3*2=-6 x^1*2=2

OpenStudy (anonymous):

\[\left(\begin{matrix}z^2 \\ x^10y^6\end{matrix}\right)\]

OpenStudy (anonymous):

now that z^0=1 \[(x ^{-5}y ^{-3})^{2}=\]?

OpenStudy (anonymous):

x^-10y^-6

OpenStudy (anonymous):

\[(x ^{b})^{a}=x ^{ab}\] multiply now the power here

OpenStudy (anonymous):

x^-5*2

OpenStudy (anonymous):

=10

OpenStudy (anonymous):

yess correct do that to y also

OpenStudy (anonymous):

y^-3*2=-6

OpenStudy (anonymous):

yess but dont forget about the sign :D

OpenStudy (anonymous):

so it's z^2/x^10y^6

OpenStudy (anonymous):

hmm z is not supposed to be there any more since its value is 1 anything you multiply by 1 is itself like ax1=a

OpenStudy (anonymous):

oh so it's 1/x^10y^6

OpenStudy (anonymous):

yesssssssssssss :D

OpenStudy (anonymous):

its right!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

YW have fun now and enjoy solving good luck ......... :D

OpenStudy (anonymous):

would u mind helping me with the other one i have on line

OpenStudy (anonymous):

ok yesss where is it? @hannalafave

OpenStudy (anonymous):

it did not work but i got one for u now 12a^3-3a^2-5a+10/4a+3

OpenStudy (anonymous):

ok the easy one to do for this one is by long division can you try that?

OpenStudy (anonymous):

yea can u go to link i sent u?

OpenStudy (anonymous):

ok